OpenStudy (anonymous):

Mrs. Sunshine has 60 folders, 45 pairs of scissors and 30 rulers. What is the greatest common factor Mrs. Sunshine can use to divide the school supplies into equal groups?

OpenStudy (whpalmer4):

You can find the greatest common factor (GCF) by factoring each number: 30 = 2*15 = 2*3*5 45 = 3*15 = 3*3*5 60 = 2*30 = 2*2*15=2*2*3*5 Now take all the numbers which appear in each number's factors and multiply them together. The product is your greatest common factor. What is the GCF of 30, 45, and 60?